how do you determine what the x and y is? the problem is {4x+2y= -2
{y = 6x-5

Answer by stanbon(75887)

determine what the x and y is?
the problem is
{4x+2y= -2
{y = 6x-5
----------------
Substitute for "y" and solve for "x":
4x + 2(6x-5) = -2
4x + 12x-10 = -2
16x = 8
x = 1/2
---------
Solve for "y":
y = 6x-5
y = 6(1/2)-5
y = 3-5
y = -2
--------------
Cheers,
Stan H.
==================